a digital camera with an original tag off 300$ was marked down by 20% you have a coupon good for additional 10% of. what is the equivalent percent discount
Answer by ikleyn(52781)

.
The equivalent percent discount is  1 - (1-0.2) * (1-0.1) = 1 - 0.8 * 0.9 = 1 - 0.72 = 0.28,   or   28%.